# What are the Essential Factors in Crafting Effective Long Context Multi-Hop Instruction Datasets? Insights and Best Practices

### v1

@@ 3.1.2 Verification Conditions
More scoring perspectives reduce long-context bias.
As illustrated in Figure 5 (a), incorporating more scoring perspectives significantly enhances the accuracy and robustness of filtering long-context data. Unlike short contexts, long contexts introduce noticeable bias in judgments. When fewer than 3 perspectives are used, performance gains are minimal, and the model often overestimates irrelevant samples, leading to poor selection results. However, increasing the number of perspectives markedly improves labeling accuracy, effectively mitigating biases associated with longer contexts. See Appendix A.2.2 for more details.
Effective verifiers adhere to annotation standards aligned with human judgment.
To assess whether incorporating additional scoring criteria enhances the model’s verification performance, we specify the criteria for each score in detail. As illustrated in Figure 5 (b), interestingly, the guideline does not include supplementary information during the annotation process for advanced models. This observation suggests that effective verifiers inherently follow annotation standards that well align with human judgment.
@@
Our methodology necessitates extension across numerous domains, emphasizing the criticality of robustness across diverse domains. Contextualizing the role of CoT (Wei et al., 2022; Qin et al., 2023), we evaluate model performance across various domains, specifically in wiki-like knowledge and paper analysis domains. As illustrated in Figure 5 (c), incorporating rationale enables the model to maintain high performance across diverse contexts. Without rationale, performance decreases by more than 8.6% when confronted with different domains. Conversely, adding rationale during validation results in minimal performance variation, with fluctuations in precision limited to 1.8% at most.
@@ 3.2.1 Generation Backbone
In practice, effective models must be capable of synthesizing high-quality data. To this end, we explored the suitability of several commonly used LLMs for single-hop data synthesis.
Figure 6: Performance of different models for generating single-hop questions.
@@
As shown in Figure 6, smaller open-source LLMs demonstrate high retention rates with cost-efficientiveness, reflecting their capability to understand and generate single-hop questions from a given context.
@@
As shown in Figure 6, more advanced LLMs increase data retention and enhance the quality of generated questions. However, these improvements are not cost-proportional, raising concerns about the economic viability of employing stronger models for single-hop question generation.
Figure 7: The analysis of generation strategies in single-hop question generation agent.
@@ 3.2.2 Generation Strategy
Furthermore, we explore whether employing a question-then-answering approach, supplemented by rationale, enhances the quality of synthetic single-hop questions.
@@
To assess whether a single or multiple stage of generation is more effective, we compare two sample generation strategies: unified question-answer and question-then-answer generation. As shown in Figure 7 (a), generating the question before the answer substantially enhances data quality. It improves both the retention rate and the data quality score, especially open-sourced LLMs, confirming its superiority. For more implementation details, see Appendix A.2.3.
@@
As illustrated in Figure 7 (b), adding rationale makes questions more relevant and insightful with higher quality. However, the improvement brought by the rationale is minimal, while the token consumption triples, making it economically inefficient.
Figure 8: The analysis of multiple question sampling.
@@ 3.3 Multiple Question Sampling
3.3.1 Retrival Strategy
This strategy involves identifying relevant documents and constructing a semantic relevance matrix to guide sampling based on both keyword and semantic scoring of documents and questions. Observations on these strategies include:
@@ 3.3.1 Retrival Strategy
We assess the effectiveness of various similarity measures by examining three metrics: embedding similarity (using BGE embeddings (Xiao et al., 2023)), BM25, and LDA. As shown in Figure 8 (a), BGE embeddings enable the model to select more relevant questions, enhancing sample quality.
@@
We also explore which aspects most influence sample quality. Figure 8 (b) demonstrates that question-based sampling significantly outperforms document-based strategies, as questions provide more contextual relevance.
@@ 3.3.2 Sampling Strategy
Figure 9: Performance Comparison of multiple question sampling based on different sampling strategies.
It selects semantically related and complementary questions from within and across documents to form coherent and contextually rich multi-hop questions.
@@
As shown in Figure 9, sampling questions within the same document results in more coherent and contextually aligned questions. However, this method may limit question diversity since they all stem from the same source.
@@
As shown in Figure 9, sampling questions from multiple documents introduces a broader range of topics and contexts, enhancing diversity. However, this increased diversity can reduce the coherence and relevance of questions due to larger topic gaps.
Figure 10: Performance of different models for merging multi-hop questions.
@@ 3.4.1 Merging Backbone
We use LLM to merge sampled questions and answers into meaningful multi-hop versions, ensuring logical consistency and contextual accuracy with the help of 5 classic LLMs. The observations are as follows:
@@
As shown in Figure 10, all models are greatly capable of handling complex question generation tasks that require multiple steps of reasoning or integration of information.
Figure 11: The analysis of multi-hop question merging agent.
@@
Figure 12: Main accuracy results by evaluation by GPT-4o, where all benchmarks comes from the LongBench (Bai et al., 2024b).
3.4.2 Merging Strategy
Question-answer pairs are enough for multi-hop instruction merging.
To minimize input tokens, we assess if long documents are necessary for enhancing merging performance. Figure 11 (a) shows that adding documents often fails to consistently improve performance and instead increases input tokens. Thus, simple question-answer pairs effectively achieve multi-hop merging.
Merging with rationale can not improve the merging quality.
Generally, generating content with rationales can enhance its quality (Qin et al., 2023; 2024). However, as depicted in Figure 11 (b), unlike single-hop generation, rationales in a multi-hop generation do not aid in forming coherent and logical questions. Our quantitative analysis further reveals that large models often misinterpret rationales within queries and merging strategies, leading to frequent CoT failures. Thus, multi-hop synthesis should avoid using additional rationales.
@@ 4 Data Utilization
4.1 Instruction Dataset Construction
To expand the domain coverage and handle longer contexts, we extended the instruction fine-tuning data across 9 domains and 2 languages. All base documents were sourced from pre-trained datasets to prevent data leakage. Our Long Multi-hop Instruction-Tuning dataset (LongMIT) results in a retention rate of over 90% in GPT-4o verification in 200 sampled samples, confirming the high quality and generalizability of our pipeline. To balance the cost and effectiveness of generating data, LongMIT are generated based on Qwen2-72B-Instruct, and verified based on InternLM2-20B. See Appendix A for more details.
@@ 4.2 Data Synthesis Efficiency
Given the high cost of data generation, we consider both cost and data quality when synthesizing LongMIT. To assess the effectiveness of this balance, we compare the proportion of high-quality data and the token cost for 200 samples generated under different strategies. As shown in Figure 13, strategies with open-source models achieve a high-quality proportion even comparable to the highest quality strategies with GPT4o, but at only one-third of the token cost. Furthermore, our approach significantly enhances data quality with minimal additional token expense compared to traditional methods. For more implementation details, see Appendix B.
4.3 Previous Instruction Dataset
(1) ChatQA (Liu et al., 2024b) uses manually annotated long text instruction-following data. (2) LongAlign (Bai et al., 2024a) leverages Claude’s generative abilities to create 10K QA pairs for training. (3) LongAlpaca (Chen et al., 2024b) integrates a large amount of paper QA corpus with additional short instruction-following examples. (4) NQ (Kwiatkowski et al., 2019) is a human-annotated long-context data with a series of natural questions.
@@ 4.4 The Results of Instruction-Tuning
Based on a substantial volume of synthesized data, we conduct instruction-tuning to further assess its utility. As shown in Table 12, our synthesized data significantly enhances the long-context QA capabilities of various LLMs, achieving an average improvement of at least 7.54% on average. Notably, multi-hop benchmarks like 2WikiMQA (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and HotpotQA (Yang et al., 2018) show more pronounced improvements. Moreover, as shown in the case study in Appendix D, the logically complex and high-quality nature of this data enables the model to generalize to single-hop QA tasks not encountered during the instruction tuning phase, further confirming the reliability of our synthetic data. Detailed procedures are available in Appendix C.
@@ 4.5 Scaling Analysis
Data Scaling Analysis
To evaluate how the size of high-quality data affects model performance, we experiment on LLaMA3-8B (Dubey et al., 2024) by varying the training data volume. The results, depicted in Figure 15, illustrate a clear relationship between the amount of data and the performance. As the dataset size increases, model performance adjusts accordingly, demonstrating the significance of high-quality data scaling in enhancing the model efficacy.
Hop Scaling Analysis
To assess the impact of multi-hop data on model performance, we increased the number of hops in the dataset while keeping the training data volume constant. This approach isolated the effect of multi-hop reasoning on model outcomes. As indicated in Figure 16, there is a clear positive correlation between the number of hops and model performance. The data demonstrate that with more hops, the model achieves higher accuracy and robustness. These results demonstrate the effectiveness of using high-quality multi-hop data to enhance the model’s capability for complex reasoning tasks.
@@ 5 Related Work
Recent efforts have aimed to enhance the performance of LLMs in handling longer contexts. LongLLaMA (Xiong et al., 2023) demonstrates the impact of incorporating long text data during various pre-training stages. LLaMA2-80K (Fu et al., 2024b) highlights the significance of using a domain-balanced, upsampled long text corpus to improve long text capabilities, requiring only a 5B-token corpus for effective comprehension. ICLM (Shi et al., 2024) enhances long-text reasoning by transforming pre-training data into knowledge graphs and splicing adjacent documents. To improve the model’s ability to follow long text instructions, LongAlpaca (Chen et al., 2024b) combines a 9K paper question-answering (QA) corpus with 3K short instruction samples. In contrast, LongAlign (Bai et al., 2024a) utilizes Claude (Anthropic, 2023) to produce 10K QA pairs for training. Additionally, ChatQA (Liu et al., 2024b) enhances long-context QA performance by incorporating manually annotated data. Building on these approaches, ChatQA2 (Xu et al., 2024a) further incorporate existing long-text datasets, such as Natural Questions (NQ) (Kwiatkowski et al., 2019).
The method closest dataset is Quest (Gao et al., 2024), which constructs QA pairs from document data and splices documents based on QA pair correlations, resulting in a close-sourced single-hop QA corpus. In contrast, our approach models document correlations first, then create multi-hop QA pairs using related intra-document data. Additionally, we offer systematic analysis, open-source datasets, and significantly improved models.
@@ 6 Conclusion
In conclusion, our proposed Multi-agent Interactive Multi-hop Generation (MIMG) framework, which includes a quality verification agent, a single-hop question generation agent, a multiple question sampling strategy, and a multi-hop question merger agent, achieves high-quality, diverse instruction data. Our experiments show that this synthetic data notably enhances performance, even surpassing models trained on larger human-annotated data, highlighting the effectiveness of our approaches.
@@
Appendix A Data Construction Details

### v2

@@ 3.1.2 Verification Conditions
More scoring criteria reduce long-context bias.
As shown in Figure 4 (a), incorporating more scoring criteria enhances the accuracy and robustness of filtering long-context data. Unlike short contexts, long contexts are prone to judgment bias. When fewer than three criteria are used, performance improvements are limited, with models often overestimating irrelevant samples. Increasing the number of criteria improves labeling accuracy, reducing biases linked to longer contexts (see Appendix C.4.2 for further details).
Effective verifiers adhere to annotation standards aligned with human judgment without guidelines.
To assess whether incorporating additional guidelines improves verification performance, we analyze the effectiveness of the method of integrating guidelines. Figure 4 (b) reveals that advanced verifiers do not require supplementary guideline information during annotation. This suggests that effective verifiers naturally adhere to annotation standards that align with human judgment.
@@
By examining CoT (Wei et al., 2022; Qin et al., 2023), in various domains like wiki knowledge and paper analysis, we observe that incorporating rationale improves model performance across diverse long-context scenarios. As shown in Figure 4 (c), without rationale, performance drops by more than 8.6% across different domains. However, adding rationale during validation minimizes performance variation, with precision fluctuations limited to 1.8%.
@@ 3.2.1 Generation Backbone
In practice, effective LLMs must be able to synthesize high-quality data. Thus, we evaluated several widely used LLMs for single-hop data synthesis.
@@
As shown in Figure 5 (a), smaller open-source LLMs exhibit high retention rates and cost-effectiveness, demonstrating their ability to generate single-hop questions from a given context.
@@
As shown in Figure 5 (a), more advanced LLMs enhance data retention and question quality. However, these improvements are not cost-proportional, raising concerns about their economic feasibility for single-hop question generation.
@@ 3.2.2 Generation Strategy
Furthermore, we explore whether a question-then-answer approach, supported by rationale, improves the quality of synthetic single-hop questions.
@@
To evaluate the effectiveness of single versus multi-stage generation, we compare two strategies: unified question-answer and question-then-answer. As shown in Figure 5 (b), generating the question first significantly improves data quality, particularly for open-sourced LLMs, increasing retention and quality scores. For implementation details, refer to Appendix C.4.3.
Figure 5: The analysis of generation backbone and generation strategies in SQGA.
Figure 6: The analysis of multiple question sampling.
@@
As illustrated in Figure 5 (c), adding rationale makes questions more relevant and insightful with higher quality. However, the improvement brought by the rationale is minimal, while the token consumption triples, making it economically inefficient.
@@ 3.3 Multiple Question Sampling
3.3.1 Retrieval Strategy
This strategy focuses on identifying relevant documents and constructing a semantic relevance matrix to guide sampling. Observations include:
@@ 3.3.1 Retrieval Strategy
We evaluate the effectiveness of various similarity measures using three metrics: embedding similarity (with BGE embeddings (Xiao et al., 2023)), BM25, and LDA. As shown in Figure 6 (a), BGE embeddings enable the model to select more relevant questions, thereby improving sample quality.
@@
We further investigate the factors affecting sample quality. Figure 6 (b) reveals that question-based sampling significantly outperforms document-based approaches, as questions provide greater contextual relevance.
Figure 7: The analysis of multi-hop question merging agent.
@@ 3.3.2 Sampling Strategy
This approach selects semantically related and complementary questions from within and across documents to create coherent, contextually rich multi-hop questions.
@@
As shown in Figure 6 (c), sampling questions within a single document yields more coherent and contextually aligned questions. However, this method may reduce diversity, as the questions are all drawn from the same source.
@@
As demonstrated in Figure 6 (c), sampling from multiple documents introduces a wider range of topics, enhancing diversity. However, this broader scope may diminish coherence and relevance due to the larger topic gaps.
@@ 3.4.1 Merging Backbone
We use LLMs to merge sampled questions and answers into coherent multi-hop versions, ensuring logical consistency and contextual accuracy with the aid of 5 classic LLMs. The key observations are as follows:
@@
As shown in Figure 7 (a), all models demonstrate strong capabilities in handling complex question-generation tasks requiring multi-step reasoning or information integration.
3.4.2 Merging Strategy
Question-answer pairs are enough for multi-hop instruction merging compared with documents.
To minimize input tokens, we examine whether long documents are necessary for improving merging performance. Figure 7 (b) shows that adding documents does not consistently enhance performance and instead increases input tokens. Thus, simple question-answer pairs are sufficient for effective multi-hop merging.
Merging with rationale can not improve the merging quality.
Although generating content with rationales generally enhances quality (Qin et al., 2023, 2024), Figure 7 (c) demonstrates that, unlike single-hop generation, rationales in multi-hop generation fail to contribute to coherent and logical question formation. Our analysis further reveals that large models often misinterpret rationales in queries and merging strategies, leading to frequent failures in CoT reasoning. Therefore, multi-hop synthesis should avoid using rationales.
@@ 3.4.2 Merging Strategy
Table 1: Main accuracy results by evaluation by GPT-4o, where all benchmarks come from the LongBench (Bai et al., 2024b). More evaluation results on Ruler (Hsieh et al., 2024) are shown in Table 4.
@@ 4.1 Data Synthesis Efficiency
Given the high cost of data generation, we balance cost and data quality in synthesizing long multi-hop instruction tuning (LongMIT) datasets (See Appendix C for more data details). To assess this balance, we compare the proportion of high-quality data and token cost for 200 samples generated under different strategies. At the input level, our interactions increase the token count by approximately 3k tokens (roughly 2.5 times improvement), which is negligible in comparison to the substantial volume of long-context documents provided, averaging 70k tokens as shown in Figure 9. At the output level, the increase in token consumption is almost no noticeable increase, with an average increase of fewer than 0.5k tokens. Despite this minimal change in token usage, the quality of our results has improved nearly fourfold. As shown in Figure 8, strategies using open-source LLMs achieve a high-quality proportion comparable to the best strategies, at only one-third of the token cost. Additionally, our approach improves data quality with minimal token expense compared to traditional methods. Then, we will pad the context with additional documents to the target length to create sufficiently long samples. For further experimental results and details, see Appendix D. Moreover, additional exploration for long code instruction data generation can be found in Appendix G.
@@ 4.2 The Results of Instruction-Tuning
We conduct instruction-tuning on the synthesized data to evaluate its utility. As shown in Table 1, our data significantly improves long-context QA capabilities across various LLMs, with an average gain of at least 7.54%. Multi-hop benchmarks such as 2WikiMQA (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and HotpotQA (Yang et al., 2018) show more notable improvements. Detailed training procedures are in Appendix E. Furthermore, in Appendix F, the high quality and logical complexity of this data enable the model to generalize to single-hop tasks not encountered during instruction tuning, confirming the reliability of MIMG.
@@ 4.3 Scaling Analysis
4.3.1 Data Scaling Analysis
To evaluate how the size of high-quality data affects model performance, we experiment on LLaMA3-8B (Dubey et al., 2024) by varying the training data volume. The results, depicted in Figure 10 in Appendix, show a clear relationship between the size of data and the performance. As the dataset size increases, model performance adjusts accordingly, demonstrating the significance of high-quality data scaling in enhancing the model efficacy.
4.3.2 Hop Scaling Analysis
To assess the impact of multi-hop data on model performance, we increased the number of hops in the dataset while keeping the training data volume constant. This approach isolated the effect of multi-hop reasoning on model outcomes. As indicated in Figure 11 in Appendix, there is a clear positive correlation between the number of hops and model performance. The data demonstrate that with more hops, the model achieves higher accuracy and robustness, demonstrating the effectiveness of using high-quality multi-hop data to enhance the model’s capability for complex reasoning tasks.
@@ 5 Related Work
Recent efforts have aimed to enhance the performance of LLMs in handling longer contexts (Hu et al., 2024; Liu et al., 2025). LongLLaMA (Xiong et al., 2023) demonstrates the impact of incorporating long text data during various pre-training stages. Then, LLaMA2-80K (Fu et al., 2024b) highlights the significance of using a domain-balanced, upsampled long text corpus to improve long text capabilities, requiring only a 5B-token corpus for effective comprehension. Furthermore, ICLM (Shi et al., 2024) enhances long-text reasoning by transforming pre-training data into knowledge graphs and splicing adjacent documents. To improve the model’s ability to follow long text instructions, LongAlpaca (Chen et al., 2024c) combines a 9K paper question-answering (QA) corpus with 3K short instruction samples. In contrast, LongAlign (Bai et al., 2024a) utilizes Claude (Anthropic, 2023) to produce 10K QA pairs for training. Additionally, ChatQA (Liu et al., 2024b) enhances long-context QA performance by incorporating manually annotated data. Building on these approaches, ChatQA2 (Xu et al., 2024a) further incorporate existing long-text datasets, such as Natural Questions (NQ) (Kwiatkowski et al., 2019).
The method closest dataset is Quest (Gao et al., 2024), which constructs QA pairs from spliced documents, resulting in a close-sourced single-hop QA corpus. In contrast, our approach models document correlations first, then create multi-hop QA pairs using related intra-document data. Additionally, we offer systematic analysis, open-source datasets, and significantly improved models.
@@ 6 Conclusion
In conclusion, our proposed Multi-agent Interactive Multi-hop Generation (MIMG) framework, including a quality verification agent, a single-hop question generation agent, a multiple question sampling strategy, and a multi-hop question merger agent, achieves high-quality, diverse instruction data. Our experiments show that this synthetic data notably enhances performance, even surpassing larger human-annotated data, highlighting the effectiveness of our approaches.
Limitations
Due to the high costs associated with large-scale distillation training experiments on GPT-4-based methods, we focus our evaluation efforts on a small-scale assessment conducted through manual evaluation. To ensure robustness, the manual annotation is performed on a sample of 200 items. While the sample size is manageable, we acknowledge the potential for minor unavoidable biases inherent in random sampling. Moreover, considering the nature of randomization for LLMs, it is quite hard for us to strictly control LLMs to generate the same question for quality comparison.
Ethical Considerations
Participants are recruited from universities across China, and all must have passed the CET-6 exam or achieved an IELTS score of 6 or higher. While participants come from diverse regions, we minimize the impact of national biases by focusing primarily on long context data. All annotators provided informed consent and were compensated above the local minimum wage. Additionally, no IRB review was required for the study.
The annotation process starts with an onboarding test, where participants answer 20 example questions. They receive \$42 for this phase, designed to familiarize them with the task. Annotators are then paid \$5 per hour, with a total of approximately 120 human hours dedicated to manual annotations. Overall, three experts are involved in the annotation and verification stages.
Acknowledgments
Thanks to Hanqi Li, Yifei Yang and Da Ma for their constructive feedbacks.
@@ Acknowledgments
Appendix A Metrics Utilized in Exploration
A.1 Metric Definition & Implementation
A.1.1 Diversity
Diversity measures the frequency of samples with the same semantics appearing in the data. In annotation, annotators sequentially identify whether new samples are semantically equivalent to previously annotated ones.
High diversity indicates a broad range of samples, ensuring that annotations do not repeat similar or identical meanings. This is essential for creating a dataset that represents various use cases and scenarios, covering a wide array of semantic topics. A diverse dataset contributes to a more robust LLM by capturing the nuances of language, context, and conceptual meaning across different long-context samples.
A.1.2 Multi-Hop
Multi-hop refers to the need for multiple information connections and integrations from various sources when handling complex samples. Here, annotation tasks require annotators to assess a query’s needs by utilizing information from several contextual documents.
Effective multi-hop datasets pose questions that cannot be answered with a single data point but demand the combination of multiple facts or steps to reach the correct answer. Such reasoning is vital in real-world long-context tasks, such as answering questions that require complex deductions or understanding interconnected pieces of information.
A.1.3 High-Quality
High-quality annotations denote the accuracy, consistency, and relevance of the synthesis data. In high-quality datasets, annotators are required to judge whether each sample is precise and reliable, minimizing errors or inconsistencies.
Annotators must ensure that the data they provide accurately reflects the meaning or intent of the task at hand. In long-context NLP, high-quality data is crucial for developing models that make accurate predictions, recognize subtle patterns, and perform effectively in real-world scenarios.
Dataset Name High-quality Diversity Multi-hop LongAlpaca (Chen et al., 2024c) 70.3 87.9 50.7 LongAlign (Bai et al., 2024a) 87.7 83.8 52.6 ChatQA2 (Xu et al., 2024a) 83.1 80.3 30.4 NQ (Kwiatkowski et al., 2019) 90.2 63.2 10.3 LongMIT (Ours) 94.8 88.2 94.8
Table 2: Results of the data quality of different instruction datasets, reporting the data quality score as the average of three independent manual evaluations.
A.2 The Impact of Different Metric
Furthermore, we conduct sampling and annotation of several instruction-tuning datasets based on three metrics. A comparison with Table 2 shows that the results require considering multiple factors and the specific needs of each task: NQ provides high-quality data, enhancing its effectiveness for in-domain tasks such as DuReader. However, its lack of diversity limits its performance in out-of-domain scenarios, which lowers its overall accuracy. In contrast, while LongAlign offers greater diversity, its quality is relatively lower, weakening its fundamental comprehension abilities. This is reflected in its significantly poorer performance on basic tasks like DuReader, resulting in suboptimal overall performance.
For the metric view, the conclusions are as follows: (1) High-quality annotations notably improve model performance in foundational long-context comprehension tasks, such as DuReader. (2) Annotation diversity is essential for enhancing model performance across a wide range of tasks. (3) The multi-hop property is particularly important for complex reasoning tasks that require integrating multiple long-context clues, as seen in datasets like HotpotQA and MusiQue.
In summary, we argue that quality ensures strong performance in fundamental long-context tasks, multi-hop promotes the complex multi-hop capabilities, while diversity improves performance across various fields by leveraging these basic capabilities.
A.3 Automatic Metrics
A.3.1 Quality Score
Several studies have shown that large-scale models align with human judgment when generating quality scores (Chen et al., 2023; Chang et al., 2024; Lee et al., 2024). These models also perform well in specialized domains, such as medicine, where they show strong consistency with expert evaluations (Haim et al., 2024). Motivated by these findings, we prompt LLMs to automatically generate quality scores for scalable and effective long-context data quality verification.
To analyze the effectiveness of quality scores in long-context scenarios, we analyze the consistency between quality scores and human assessments. Specifically, Figure 3 (a) demonstrates that the Kappa coefficient for the agreement between the scoring mechanism and human evaluators exceeds 0.50, significantly surpassing the performance of classification strategies that directly label items as high or low quality. Furthermore, Figure 3 (b) shows that the scoring mechanism achieves a precision level of 96.43% relative to human evaluators, underscoring its effectiveness as a robust data screening strategy.
A.3.2 Retention Rate
Retention rate refers to the proportion of data retained after being filtered by a quality verification agent based on a specified threshold. This concept arises from our observation that LLMs excel more as selectors than as annotators in Sec. 3.1.1, demonstrating higher retention rates and better alignment with human judgment. In complex or subtle situations, LLM judgments may vary, which can impact data interpretability and reliability. Therefore, we select models and strategies that closely align with human consistency, which achieves an almost perfect precision rate of 96.43%.
For implementation, we use a threshold of 8.5, determined through an internally annotated verification set of 200 items. The threshold yielding the highest precision in this set was chosen as the standard parameter for future evaluations.
Regarding the “average score”, our strategy assigns a quality score from 0 to 10 for each data point. After evaluating the entire dataset, we calculate the average score by averaging all individual quality scores.
Figure 9: The analysis of constructed dataset distribution.
Appendix B Discussion of Scoring versus Classification
Although the document inputs are relatively long, averaging approximately 3K tokens, the classifier still assigns a “high-quality” label to nearly every sample of this length, even when we introduce multiple reference criteria. We believe that, once inputs exceed this length threshold, the classifier’s baseline score effectively rises above five, so any score over five is automatically deemed high quality. As a result, the classifier’s judgments lose discriminative power: every long sample is labeled “high-quality”.
By contrast, the scoring approach provides a more nuanced ranking. Even if all samples are of generally high quality, a score of 0.9 is still recognized as slightly better than 0.8. This finer distinction more closely reflects human evaluation.
Appendix C Data Construction Details
C.1 Dataset Construction Pipeline Discussion
The current data in this field relies on Self-Instruct but lacks a systematic framework. Notably, many existing long-text works are essentially subsets of our MIMG. For instance, the prompts of Self-Instruction and those of the Single-hop Question Generation Agent are nearly identical. Additionally, Self-Instruction (+ LLM recheck) aligns with the architecture integrated into our Quality Verification Agent. Therefore, we assert that our approach offers a more thorough solution.